OpenStudy (anonymous):

If log_sinx(cosx)=1/2, and 0

hartnn (hartnn):

lets apply the definition, \(\log_ab=c \implies b=a^c\) you knew this, right ?

hartnn (hartnn):

what will log_sinx(cosx)=1/2 simplify to ?

OpenStudy (anonymous):

Sinx^1/2=cosx?

hartnn (hartnn):

yup, now square both sides

OpenStudy (anonymous):

Sinx= (cosx)^2

hartnn (hartnn):

try to bring everything in sin x

hartnn (hartnn):

how would u write cos^2 x in terms of sin x ?

OpenStudy (anonymous):

cos( 2x) = 1 - sin^2( x)

hartnn (hartnn):

you meant cos^2x = 1-sin^2 x then its correct :) so sin x = 1-sin^2 x can you now solve for sin x ? its just a quadratic

hartnn (hartnn):

still here ? stuck or trying ?

OpenStudy (anonymous):

Sin^2x+sinx -1=0 is this right?

hartnn (hartnn):

yup. its like y^2 + y-1 =0 where y= sin x a simple quadratic equation, solve for y, you would have solved for sin x